OpenStudy (anonymous):

If the parallel sides of a trapezoid contained by the lines y = -x + 5 and y = -x - 9, what equation represents the line contained by the midsegment?

OpenStudy (cwrw238):

well the slope of that line will be the same as the parallel lines because it will also be parallel to them so that y = -x +/- something

OpenStudy (cwrw238):

the intercept on the y axis will be halfway between these lines it = (+5-9) / 2 = -4/2 = -2 so now you've got your equation

OpenStudy (anonymous):

Thank you @cwrw238 i got y=-x-2

OpenStudy (cwrw238):

thats it yw